OpenStudy (anonymous):

a lawn company is creating a new lawn by putting down pallets of sod the crew can put down 1,200 square feet of sod in 20 minutes one pallet= 400 square feet of sod how many pallets of sod per hour can the crew put down

OpenStudy (anonymous):

Every hour they can put down 9 pallets an hour

OpenStudy (anonymous):

1200 \[1200\div400=3 \] \[60\div20=3\] 3x3=9

OpenStudy (gw2011):

Since 1200 sq ft can be done in 20 minutes, then in 60 minutes (1 hour) 3600 sq ft can be done. You can set up a proportion kto get the following answer: 1/400=x/3600 400x=3600 x=9 this is the answer

OpenStudy (anonymous):

thanks